# AgentSociety: Large-Scale Simulation of LLM-Driven Generative Agents Advances Understanding of Human Behaviors and Society

### v1

@@ Abstract
Understanding human behavior and society is a central focus in social sciences, with the rise of generative social science marking a significant paradigmatic shift. By leveraging bottom-up simulations, it replaces costly and logistically challenging traditional experiments with scalable, replicable, and systematic computational approaches for studying complex social dynamics. Recent advances in large language models (LLMs) have further transformed this research paradigm, enabling the creation of human-like generative social agents and realistic simulacra of society. In this paper, we propose AgentSociety, a large-scale social simulator that integrates LLM-driven agents, a realistic societal environment, and a powerful large-scale simulation engine. Based on the proposed simulator, we generate social lives for over 10k agents, simulating their 5 million interactions both among agents and between agents and their environment. Furthermore, we explore the potential of AgentSociety as a testbed for computational social experiments, focusing on four key social issues: polarization, the spread of inflammatory messages, the effects of universal basic income policies, and the impact of external shocks such as hurricanes. These four issues serve as valuable cases for assessing AgentSociety’s support for typical research methods – such as surveys, interviews, and interventions – as well as for investigating the patterns, causes, and underlying mechanisms of social issues. The alignment between AgentSociety’s outcomes and real-world experimental results not only demonstrates its ability to capture human behaviors and their underlying mechanisms, but also underscores its potential as an important platform for social scientists and policymakers.
@@ 1 Introduction
Indeed, the authenticity of these simulations depends on to what extent the most basic components – i.e., the agents – behave like humans. However, most existing agents, driven by rules [29], equations [45], or even machine learning models [115], are limited in their ability to generate human-like behaviors. For example, when simulating opinion dynamics, people’s opinions are often represented as scalars or vectors, and their interactions as equations [15]. While this modeling approach offers valuable insights, it is still far from reality, as people typically communicate using natural language, rather than numeric values. Fortunately, recent advances in large language models (LLMs) have shown promise in creating human-like agents [38, 99, 106]. Numerous studies have pointed out that after being empowered by LLMs, these agents agents have generated human-like “minds” [92, 63, 55]. They not only possess basic cognition abilities, such as learning [106, 99], reasoning [102], and decision-making [64, 38], but also demonstrate the capability to understand and predict the thoughts and intentions of others [55, 92]. Furthermore, beyond exploring these agents’ minds, some researchers have investigated their potential to mimic human behaviors [64, 80, 89, 108, 33, 48, 39, 77]. Their investigations have revealed that, through elaborate designs incorporating domain knowledge, these LLM-driven agents can generate social behaviors, such as mobility [89, 108, 33], employment [64, 48], consumption [64, 48], and social interactions [39, 77]. While great efforts have been made to examine specific facets of these agents, simulating a comprehensive social being remains largely underexplored.
@@
To address the above gaps, we propose AgentSociety, a large-scale social generative simulator that incorporates LLM-driven social generative agents, a realistic societal environment, and large-scale interactions both among agents and between agents and the environment. Specifically, following social theories from a broad range of fields, including psychology [68, 5], economics [25] and behavioral sciencel [118], we first design a framework for LLM-driven social agents. These agents are endowed with human-like “minds”, which include emotions, needs, motivations, and cognition of the external world. Their behaviors such as mobility, employment, consumption, and social interactions are dynamically driven by these internal mental states. Beyond individual agents, we construct a realistic societal environment that seamlessly integrates urban, social, and economic spaces, providing a rich foundation for agent interactions and self-evolution. At its core, society emerges from the bottom-up interactions among individuals, where agent-level interactions collectively give rise to complex social structures and phenomena. Recognizing that social systems exhibit emergent behaviors shaped by scale, we develop a large-scale social simulation engine equipped with distributed computing and an MQTT-powered high-performance messaging system. This enables simulations with up to 10k agents, each engaging in an average of 500 interactions per day, capturing the intricate dynamics of large-scale social systems. Based on the proposed large-scale social simulator, we successfully reproduce behaviors, outcomes, and patterns observed in four real-world social experiments, including polarization, inflammatory message spread, the effects of universal basic income policies, and the impact of external shocks like hurricanes. These experiments not only cover social research methods, such as surveys, interviews, and interventions, but also demonstrate the simulator’s ability to replicate social dynamics, unlocking new possibilities for social scientists and policymakers. Overall, AgentSociety marks a paradigm shift in AI for social science, enabling large-scale, high-fidelity simulations that overcome traditional experimental limitations in costs, scalability, and feasibility. By leveraging LLM-driven social generative agents, it facilitates deeper analysis, prediction, and intervention in complex social systems, laying the foundation for computational social science 2.0.
@@ 2 AgentSociety: Design and Overview
In this paper, we propose AgentSociety, a comprehensive large-scale social simulator designed to integrate LLM-driven social generative agents, a realistic societal environment, and a robust simulation engine. This simulator not only supports large-scale agents and their interactions, but also facilitates advanced social science research. Figure 2 provides an overview of AgentSociety and outlines the structure of this paper. AgentSociety consists of three key components: LLM-driven social generative agents, a realistic societal environment, and a powerful simulation engine that supports large-scale interactions. Extensive experiments demonstrate AgentSociety’s superior performance and its potential as a valuable testbed for various social experiments. In particular, we first introduce LLM-driven social generative agents in Section 3, which discusses the designs for agents’ minds, complex social behaviors, and their coupling in detail. We then demonstrate our real-world societal environment in Section 4, which includes our modeling of urban, social, and economic spaces. Furthermore, we illustrate our large-scale social simulation engine in Section 5 and evaluate its performance in Section 6. Finally, we show a typical one day life of our simulated agents in Section 7.1 and launch several social experiments based on our proposed large-scale social simulator in Sections 7.2 - 7.5. These examples, covering polarization (Section 7.2), the spread of inflammatory messages (Section 7.3), universal basic income 7.4, and external shocks of hurricanes (Section 7.5), demonstrate the validity and authenticity of our proposed simulator.
@@ 6 Performance Evaluation
All experiments were conducted on Huawei Cloud c7.16xlarge.4 cloud servers to ensure comparability of results. To mitigate potential interference from rate-limiting effects inherent in LLM API calls during large-scale social simulator execution, we chosen the DeepSeek API platformhttps://platform.deepseek.com/ that officially claims no request limithttps://api-docs.deepseek.com/quick_start/rate_limit. Related experiments were specifically scheduled during DeepSeek’s off-peak hours (05:00-07:00 local time) to maximize the LLM API throughput. According to a DeepSeek website statement, the model used during the experiments was DeepSeek-V3 [65].
@@ 7.1 One Day Life
Based on the Social Agent’s capability to simulate a one-day life, we further conducted simulation experiments in the domains of cognition, social interaction, economics, and mobility. These experiments were designed to validate the Social Agent’s proficiency in capturing behaviors across various domains, as illustrated in Fig.14.

### v2

@@ Abstract
Understanding human behavior and society is a central focus in social sciences, with the rise of generative social science marking a significant paradigmatic shift. By leveraging bottom-up simulations, it replaces costly and logistically challenging traditional experiments with scalable, replicable, and systematic computational approaches for studying complex social dynamics. Recent advances in large language models (LLMs) have further transformed this research paradigm, enabling the creation of human-like generative social agents and realistic simulacra of society. In this paper, we propose AgentSociety, a large-scale social simulator that integrates LLM-driven agents, a realistic societal environment, and a powerful large-scale simulation engine. Based on the proposed simulator, we generate social lives for over 10k agents, simulating their 5 million interactions both among agents and between agents and their environment. Furthermore, we explore the potential of AgentSociety as a testbed for computational social experiments, focusing on five key social issues: polarization, the spread of inflammatory messages, the effects of universal basic income policies, the impact of external shocks such as hurricanes, and urban sustainability. These five issues serve as valuable cases for assessing AgentSociety’s support for typical research methods – such as surveys, interviews, and interventions – as well as for investigating the patterns, causes, and underlying mechanisms of social issues. The alignment between AgentSociety’s outcomes and real-world experimental results not only demonstrates its ability to capture human behaviors and their underlying mechanisms, but also underscores its potential as an important platform for social scientists and policymakers.
@@ 1 Introduction
Indeed, the authenticity of these simulations depends on to what extent the most basic components – i.e., the agents – behave like humans. However, most existing agents, driven by rules [30], equations [47], or even machine learning models [119], are limited in their ability to generate human-like behaviors. For example, when simulating opinion dynamics, people’s opinions are often represented as scalars or vectors, and their interactions as equations [15]. While this modeling approach offers valuable insights, it is still far from reality, as people typically communicate using natural language, rather than numeric values. Fortunately, recent advances in large language models (LLMs) have shown promise in creating human-like agents [39, 103, 110]. Numerous studies have pointed out that after being empowered by LLMs, these agents have generated human-like “minds” [96, 66, 58]. They not only possess basic cognition abilities, such as learning [110, 103], reasoning [106], and decision-making [67, 39], but also demonstrate the capability to understand and predict the thoughts and intentions of others [58, 96]. Furthermore, beyond exploring these agents’ minds, some researchers have investigated their potential to mimic human behaviors [67, 83, 93, 112, 34, 51, 40, 80]. Their investigations have revealed that, through elaborate designs incorporating domain knowledge, these LLM-driven agents can generate social behaviors, such as mobility [93, 112, 34], employment [67, 51], consumption [67, 51], and social interactions [40, 80]. While great efforts have been made to examine specific facets of these agents, simulating a comprehensive social being remains largely underexplored.
@@
To address the above gaps, we propose AgentSociety, a large-scale social generative simulator that incorporates LLM-driven social generative agents, a realistic societal environment, and large-scale interactions both among agents and between agents and the environment. Specifically, following social theories from a broad range of fields, including psychology [71, 5], economics [25] and behavioral sciencel [122], we first design a framework for LLM-driven social agents. These agents are endowed with human-like “minds”, which include emotions, needs, motivations, and cognition of the external world. Their behaviors such as mobility, employment, consumption, and social interactions are dynamically driven by these internal mental states. Beyond individual agents, we construct a realistic societal environment that seamlessly integrates urban, social, and economic spaces, providing a rich foundation for agent interactions and self-evolution. At its core, society emerges from the bottom-up interactions among individuals, where agent-level interactions collectively give rise to complex social structures and phenomena. Recognizing that social systems exhibit emergent behaviors shaped by scale, we develop a large-scale social simulation engine equipped with distributed computing and an MQTT-powered high-performance messaging system. This enables simulations with up to 10k agents, each engaging in an average of 500 interactions per day, capturing the intricate dynamics of large-scale social systems. Based on the proposed large-scale social simulator, we successfully reproduce behaviors, outcomes, and patterns observed in five real-world social experiments, including polarization, inflammatory message spread, the effects of universal basic income policies, the impact of external shocks like hurricanes, and urban sustainability. These experiments not only cover social research methods, such as surveys, interviews, and interventions, but also demonstrate the simulator’s ability to replicate social dynamics, unlocking new possibilities for social scientists and policymakers. Overall, AgentSociety marks a paradigm shift in AI for social science, enabling large-scale, high-fidelity simulations that overcome traditional experimental limitations in costs, scalability, and feasibility. By leveraging LLM-driven social generative agents, it facilitates deeper analysis, prediction, and intervention in complex social systems, laying the foundation for computational social science 2.0.
@@ 2 AgentSociety: Design and Overview
In this paper, we propose AgentSociety, a comprehensive large-scale social simulator designed to integrate LLM-driven social generative agents, a realistic societal environment, and a robust simulation engine. This simulator not only supports large-scale agents and their interactions, but also facilitates advanced social science research. Figure 2 provides an overview of AgentSociety and outlines the structure of this paper. AgentSociety consists of three key components: LLM-driven social generative agents, a realistic societal environment, and a powerful simulation engine that supports large-scale interactions. Extensive experiments demonstrate AgentSociety’s superior performance and its potential as a valuable testbed for various social experiments. In particular, we first introduce LLM-driven social generative agents in Section 3, which discusses the designs for agents’ minds, complex social behaviors, and their coupling in detail. We then demonstrate our real-world societal environment in Section 4, which includes our modeling of urban, social, and economic spaces. Furthermore, we illustrate our large-scale social simulation engine in Section 5 and evaluate its performance in Section 6. Finally, we show a typical one day life of our simulated agents in Section 7.1 and launch several social experiments based on our proposed large-scale social simulator in Sections 7.2 - 7.5. These examples, covering polarization (Section 7.2), the spread of inflammatory messages (Section 7.3), universal basic income 7.4, external shocks of hurricanes (Section 7.5), and urban sustainability(Section 7.6), demonstrate the validity and authenticity of our proposed simulator.
@@ 6 Performance Evaluation
All experiments were conducted on Huawei Cloud c7.16xlarge.4 cloud servers to ensure comparability of results. To mitigate potential interference from rate-limiting effects inherent in LLM API calls during large-scale social simulator execution, we have chosen the DeepSeek API platformhttps://platform.deepseek.com/ that officially claims no request limithttps://api-docs.deepseek.com/quick_start/rate_limit. Related experiments were specifically scheduled during DeepSeek’s off-peak hours (05:00-07:00 local time) to maximize the LLM API throughput. According to a DeepSeek website statement, the model used during the experiments was DeepSeek-V3 [68].
@@ 7.1 One Day Life
Based on the Social Agent’s capability to simulate a one-day life, we further conducted simulation experiments in the domains of cognition, social interaction, economics, and mobility. These experiments are designed to validate the Social Agent’s proficiency in capturing behaviors across various domains as illustrated in Fig.14.
@@ 7.5 External Shocks of Hurricane
7.6 Urban Sustainability
Pro-environmental interventions in urban systems constitute a critical area of study, given their profound implications for sustainable development and collective behavioral change. Understanding how different normative strategies shape individual decision-making and aggregate mobility patterns is essential for designing effective policies that promote low-carbon lifestyles and mitigate environmental impact.
We construct a high-fidelity simulation of an urban living environment in Beijing. The simulated environment is grounded in real-world GIS data and incorporates detailed Areas of Interest (AOIs) and Points of Interest (POIs) spanning residential communities, commercial districts, office zones, and transportation hubs. Within this digital twin, we instantiate a population of 200 heterogeneous agents whose demographic attributes, including age, gender, education, and marital status, are sampled according to the 2024 Beijing census, thereby ensuring basic demographic realism and representativeness. Using this urban sandbox, we invite six independent research teams to freely design and deploy distinct eco-normative systems aimed at encouraging low-carbon travel behavior.
The experiment lasts for two simulated days. On the first day, each research team is allowed to inject its enhanced eco-normative system into the city through three intervention channels. The first channel is direct communication with selected simulated residents. This form of interpersonal outreach does not incur monetary cost, but the number of interactions is capped. The second channel is poster placement in designated AOIs, such that any simulated agent passing through these locations can observe the posted normative messages. The third channel is a city-wide public announcement, through which the research team can broadcast relevant information to all agents in the simulated city. Poster placement and city-wide announcements incur fixed costs of 1,000 and 10,000 per intervention, respectively. Each team operates under a total budget ceiling of 100,000 and is therefore incentivized to achieve the strongest possible intervention effect with minimal resource expenditure, rather than relying on unrestricted intervention intensity.
No further intervention is allowed on the second day. Instead, the effects of the injected normative systems are evaluated through simulation-based observation along two complementary dimensions. First, we measure the internalized strength of agents’ pro-environmental norm systems using a survey instrument (Table 7). The survey assesses several dimensions of environmental norm strength, including low-carbon travel awareness and transport-choice norms, household energy use and conservation norms, green consumption and sustainable purchasing tendencies, low-carbon dietary and lifestyle choices, awareness of resource conservation and recycling, and the degree of environmental responsibility and norm internalization in everyday behavior. Second, we evaluate revealed behavioral outcomes by recording the travel modes selected by simulated agents on the second day. Agents may choose either private-car travel or low-carbon alternatives such as walking, cycling, buses, and metro. Based on realized travel distance and travel mode, we further estimate the carbon emissions associated with each agent’s mobility behavior, thereby providing a behavioral and outcome-based measure of intervention effectiveness.
Figure 24: Effects of six eco-normative intervention systems on agents’ pro-environmental alignment and travel behavior. (a) Increase in pro-environmental norm strength relative to the baseline, measured by post-intervention survey responses. (b) Reduction in mobility-related CO2 emissions relative to the baseline, estimated from second-day travel distance and mode choice. (c) Relationship between the increase in pro-environmental norm strength and the reduction in CO2 emissions across the six research teams. The positive trend indicates that stronger normative internalization is associated with larger behavioral shifts toward low-carbon mobility.
Results show that all six injected normative systems produce measurable improvements at both the perceptual and behavioral levels. As shown in Figure 24a-b, all six teams increase agents’ pro-environmental normative alignment and, at the same time, reduce mobility-related CO2 emissions relative to the baseline. This suggests that the injected interventions do not merely change stated attitudes, but also translate into more frequent adoption of low-carbon travel modes in the simulated city. Moreover, Figure 24c reveals a clear positive association between the increase in pro-environmental norm strength and the reduction in CO2 emissions: teams that more successfully foster normative alignment also tend to induce larger behavioral shifts toward walking, cycling, and public transit. Taken together, these results indicate that pro-environmental norm formation is not only an attitudinal outcome in itself, but also an important mechanism through which downstream behavioral change is realized.
Figure 25: Normative composition of the six intervention systems designed by independent research teams. Each pie chart shows the proportion of four normative components in a team’s injected eco-normative system: descriptive norms, injunctive norms, personal norms, and identity norms.
To further understand why some interventions are more effective than others, we analyze the normative composition of the six submitted systems (Figure 25). Across teams, four major normative components emerge: descriptive norms, which emphasize what others commonly do; injunctive norms, which emphasize what people are expected or encouraged to do [26]; personal norms, which stress internalized moral obligation [92]; and identity norms, which connect behavior to one’s self-concept as a responsible or environmentally conscious individual [42]. A comparison across teams suggests that interventions with a stronger emphasis on personal norms are generally more effective than those relying primarily on injunctive norms. In particular, Team 3, which assigns the largest share to personal norms and a substantial share to identity norms, achieves the best overall performance on both normative alignment and emission reduction. By contrast, Team 2, whose system is dominated by injunctive norms, produces only moderate improvements. This pattern is consistent with prior evidence showing that personal norms are the strongest predictor of pro-environmental behavior and mediate much of the influence of social norms [48]. Overall, these findings suggest that interventions are more effective when they move beyond compliance-oriented messaging and instead cultivate self-endorsed moral commitment and identity-consistent low-carbon action.
Table 7: Pro-environmental awareness and norms survey.
Q# Question and options
Q1 Question: You are financially well-off and own a private car. You are going to a restaurant 3 km away. Which transportation mode would you choose? A. Walk B. Ride a bicycle or shared bike C. Take public transportation D. Drive a private car
Q2 Question: You are working alone at home. It is summer and the outdoor temperature is 35°C. For indoor cooling, what would you choose? A. Close curtains and use an electric fan B. Set AC at 26–28°C and turn it off periodically C. Keep the AC on all day for comfort D. Set the AC below 20°C to ensure maximum coolness
Q3 Question: You plan to shop at a supermarket. Regarding shopping bag use, what do you usually do? A. Bring reusable eco-friendly shopping bags B. Try to minimize bag use; use store-provided bags only when necessary C. Use a suitable number of store-provided bags according to purchase volume D. Use separate bags for each category of goods for convenience
Q4 Question: When choosing your daily diet, which description best matches your habit? A. Mostly choose seasonal fruits/vegetables and rarely eat meat B. Value dietary diversity while controlling meat intake C. Prefer meat but still pair it with some vegetables D. Prefer a high-protein diet with meat as the main food source
Q5 Question: Which of the following best describes your daily water-use behavior? A. Reuse vegetable-washing water (e.g., for plants) and keep showers within 5 minutes B. Turn off taps promptly and keep showers around 10 minutes C. Occasionally forget to fully turn off taps and prefer long comfortable showers D. Usually fill a bathtub because water cost is not a major concern
Q6 Question: When handling household waste, what do you usually do? A. Strictly sort waste into recyclable, kitchen, hazardous, and residual categories B. Mostly sort waste into recyclable vs. non-recyclable C. Know sorting is recommended but often skip it due to inconvenience D. Believe waste will be processed anyway, so sorting has little impact
Q7 Question: You work in the city center and live 6 km from your office. What is your preferred daily commuting mode? A. Bicycle or e-bike B. Public transportation C. Ride-hailing or taxi D. Private car for convenience and privacy
Q8 Question: Regarding home appliance usage, what do you usually do? A. Turn off and unplug appliances when not in use; purchase energy-efficient devices B. Use smart plugs to manage standby appliances and reduce energy use C. Turn off long-idle appliances but keep them plugged in D. Keep most appliances on standby for convenience
Q9 Question: When planning weekend leisure activities, which are you most likely to choose? A. Activities in nearby parks or community centers B. Public transportation to urban cultural venues or malls C. Drive to suburban attractions for a day trip D. Plan short-haul flights or long-distance self-driving trips
Q10 Question: Which description best matches your personal consumption habits? A. Prioritize durability and repair old items instead of buying new ones B. Consider necessity before shopping; occasionally pay more for quality C. Follow trends and update items, often buying during promotions D. Prefer the latest products and frequently replace phones, clothing, etc.